Why is the answer B? I keep getting C (using old wage x new cpi divided by old cpi)

39. President von Clownstick and former President Vacuum are having an argument President von Clownstick says that he is not paid enough, and that he wishes he had been in office when President Vacuum was. President von Clownstick earns $400,000 per year, increased from 12.6 when Vacuum left office to 246.5 today. Who is earning more in real terms, and what would both presidents salaries be in the base year? while President Vacuum earns $50,000 per year. At the same time, CPI has A President von Clownstick is better off; President von Clownsticks salary would be B President Vacuum is better off; President von Clownsticks salary would be $162,271 C President Vacuum is better off; President von Clownsticks salary would be $20,466 D Impossible to know without knowing the Ruble-USD exchange rate $7,825,396 and President Vacuums would be $2,555 per year and President Vacuums would be $396,825 and President Vacuums would be $978,174
Answer #1
president Nominal Salary CPI Real Salary
Vaccum 50,000 12.6 50000/12.6= 3968.25
Clownstick 400,000 246.5 400,000/246.5 = 1622.71

Thus salary in real terms is nominal Salary/ CPI

So in real terms, salary of vaccum is higher

Thus president vaccum is better off,

Now assuming base year CPI equal to 100, then nominal Salary in base year is real salary * CPI

= 1622.71*100 = 162,271 : clownstick salary

= 3968.25*100 = 396,825: vaccum salary

Thus answer is option b)